# Online Multi-modal Root Cause Analysis

### v1

@@ Appendix D Evaluation Metrics
We evaluate the model performance with the following four widely-used metrics Meng et al. (2020):
@@
(1). Precision@K (PR@K): It measures the probability that the top $K$ predicted root causes are relevant by:
$PR@K=\frac{1}{|\mathbb{A}|}\sum_{a\in\mathbb{A}}\frac{\sum_{i<k}R_{a}(i)\in V_{a}}{\min(K,|v_{a}|)}$ (22)
@@
(2). Mean Average Precision@K (MAP@K): It evaluates the top $K$ predicted causes from the overall perspective formulated as:
$MAP@K=\frac{1}{K|\mathbb{A}|}\sum_{a\in\mathbb{A}}\sum_{i\leq j\leq K}PR@j$ (23)
@@
(3). Mean Reciprocal Rank (MRR): It assesses the ranking capability of models, defined as:
$PR@K=\frac{1}{|\mathbb{A}|}\sum_{a\in\mathbb{A}}\frac{1}{rank_{R_{a}}}$ (24)
@@
where $rank_{R_{a}}$ is the rank number of the first correctly predicted root cause for system fault $a$.
(4). Time: Measures the training time (in seconds) for each batch of data.
Appendix E Additional Experiment
E.1 Experimental Results
We report the experimental results on Train Ticket dataset in Table 5.
Table 5: Results on Train Ticket w.r.t different metrics.
Modality Model PR@1 PR@5 PR@10 MRR MAP@3 MAP@5 MAP@10 Time (s)
Metric Only PC 0 0 0.2 0.067 0 0 0.06 9.65
Dynotears 0 0 0 0.047 0 0 0 21.3
C-LSTM 0 0.2 0.2 0.097 0 0.08 0.14 30.63
GOLEM 0 0.2 0.2 0.098 0 0.08 0.14 9.56
REASON 0.2 0.4 0.6 0.323 0.2 0.36 0.48 9.03
CORAL 0 0.4 1.0 0.184 0 0.16 0.5 5.72
Log Only PC 0 0.2 0.4 0.166 0.133 0.16 0.26 6.34
Dynotears 0 0 0.2 0.072 0 0 0.02 17.26
C-LSTM 0 0 0.2 0.072 0 0 0.02 21.40
GOLEM 0 0.2 0.6 0.125 0 0.08 0.24 7.54
REASON 0 0.2 0.6 0.126 0 0.08 0.28 7.58
CORAL 0 0.2 0.8 0.138 0 0.08 0.32 3.35
Multi-Modality PC 0 0 0.2 0.083 0 0 0.1 12.83
Dynotears 0 0.4 0.6 0.141 0 0.16 0.32 27.82
C-LSTM 0.2 0.4 0.6 0.294 0.2 0.28 0.36 36.76
GOLEM 0 0.4 0.6 0.144 0 0.16 0.3 12.16
REASON 0.2 0.4 0.6 0.300 0.2 0.28 0.42 12.81
MULAN 0.2 0.4 1.0 0.317 0.2 0.28 0.46 11.42
CORAL 0.2 0.4 1.0 0.334 0.2 0.28 0.56 7.26
OCEAN 0.2 0.4 1.0 0.381 0.333 0.36 0.58 3.22
E.2 Additional Ablation Study
We conduct additional ablation study by replacing the dilated convolutional neural network with LSTM or Transformer model on the Product Review and Train Ticket datasets in Table 6. The results showed a performance decline, demonstrating the effectiveness of our design.
Table 6: Additional Ablation Study
@@ E.2 Additional Ablation Study
E.3 Additional Parameter Analysis
@@
In this subsection, we conduct a comprehensive parameter sensitivity analysis on the Online Boutique and Train Ticket datasets. Similarly, we assess the impact of three parameters on the overall objective functions (Eq. 3.4): $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$. Figures 3 and Figures 4 show the experimental results in terms of Mean Reciprocal Rank (MRR) on the Online Boutique and Train Ticket datasets. The x-axis represents $\ln(\lambda_{i})$, where $i\in[1,2,3]$, and the y-axis indicates the MRR score. We consistently observe that OCEAN favors a larger value of $\lambda_{1}$ on these two datasets as the temporal causal structure learning module is crucial in capturing both temporal and causal dependency. Different from the parameter analysis on the Product Review dataset, we find that $\lambda_{2}$ and $\lambda_{3}$ are not very sensitive on the Online Boutique and Train Ticket datasets. We conjecture that this can be attributed to the small size of these two datasets and both sparse regularization and acyclic constraint contribute less to securing high performance on these two datasets than the Product Review dataset.
Appendix F Reproducibility
All experiments are conducted on a desktop running Ubuntu 18.04.5 with an Intel(R) Xeon(R) Silver 4110 CPU @2.10GHz and one 11GB GTX2080 GPU. In the experiment, we set the size of historical metric and log data to 8-hour intervals and each batch is set to be a 10-minute interval. We use the Adam as the optimizer and we train the model for 100 iterations at each batch. We use two layers of dilated convolutional operations in the experiment. As for the stopping criteria, we terminate the identification process if the similarity $\gamma$ between the current batch and the previous batch is greater than 0.9 for three consecutive times.

### v2

@@ A-A Evaluation Metrics
We evaluate the model performance with the following four widely-used metrics [wang2023interdependent, DBLP:conf/iwqos/MengZSZHZJWP20]:
@@
(1) Precision@K (PR@K): It measures the probability that the top $K$ predicted root causes are relevant by:
$\text{PR@K}=\frac{1}{|\mathbb{A}|}\sum_{a\in\mathbb{A}}\frac{\sum_{i<k}R_{a}(i)\in V_{a}}{\min(K,|v_{a}|)}$
@@
(2) Mean Average Precision@K (MAP@K): It evaluates the top $K$ predicted causes from the overall perspective formulated as:
$\text{MAP@K}=\frac{1}{K|\mathbb{A}|}\sum_{a\in\mathbb{A}}\sum_{i\leq j\leq K}PR@j$
@@
(3) Mean Reciprocal Rank (MRR): It assesses the ranking capability of models, defined as:
$\text{MRR@K}=\frac{1}{|\mathbb{A}|}\sum_{a\in\mathbb{A}}\frac{1}{\text{rank}_{R_{a}}}$
@@
where $\text{rank}_{R_{a}}$ is the rank number of the first correctly predicted root cause for system fault $a$.
(4) Time: Measures the training time (in seconds) for each batch of data.
A-B Reproducibility
All experiments are conducted on a desktop running Ubuntu 18.04.5 with an Intel(R) Xeon(R) Silver 4110 CPU @2.10GHz and one 11GB GTX2080 GPU. In the experiment, we set the size of historical metric and log data to 8-hour intervals and each batch is set to be a 10-minute interval. We use the Adam as the optimizer and we train the model for 100 iterations at each batch. We use two layers of dilated convolutional operations in the experiment. As for the stopping criteria, we terminate the identification process if the similarity $\gamma$ between the current batch and the previous batch is greater than 0.9 for three consecutive times. Each experiment is run with one trial.
A-C Time Complexity
The time complexity of dilated convolution based causal structure learning is $O(nTkC)$. Here, $n$ is the number of system entities, $T$ is the sequence length, $L$ is the number of layers, $C$ is the number of filters, and $k$ is the filter size. The time complexity of multi-factor attention module is $O(d_{M}d_{L}d)$, where $d$ is the hidden feature dimensionality, and $d_{M}$ and $d_{L}$ are the number of factors in two modalities. The time complexity of contrastive learning module is $O(n^{2}d)$. Although Transformers (or LLMs) could capture temporal dependencies, they often come with high computational overhead. Specifically, the complexity of the Transformer is $O(nLT^{2}d)$. Notice that $T$ can be very large on some large-scale datasets, such as Product Review dataset [zheng2024lemma]. As a result, Transformers are computationally more expensive than dilated convolutional neural networks. We further demonstrate the efficiency and effectiveness of our method compared to LSTM and Transformers in the ablation study presented in Table VII.
A-D Baselines
We compare OCEAN with eight causal discovery based RCA methods: (1) PC [DBLP:journals/technometrics/Burr03]: A classic constraint-based algorithm that identifies the causal graph’s skeleton using an independence test. (2) Dynotears [DBLP:conf/aistats/PamfilSDPGBA20]: A dynamic Bayesian network that uses vector autoregression model to capture the temporal dependency and learn the acyclic graph with DAG constraint. (3) C-LSTM [DBLP:journals/pami/TankCFSF22]: A root cause analysis method using LSTM to model temporal dependencies and capture nonlinear Granger causality. (4) REASON [wang2023interdependent]: A causal structure learning method that Learns both intra-level and inter-level causal relationships in interdependent networks. (5) MULAN [DBLP:conf/www/ZhengCHC24]: A multi-modal method that captures both modality-invariant and modality-specific representations. (6) $\epsilon$-Diagnosis [shan2019diagnosis]: A lightweight, unsupervised algorithm that detects root causes of long-tail latency in microservices by comparing time series similarity using $\epsilon$-statistics. (7) BARO [pham2024baro]: An end-to-end RCA framework that combines change point detection and robust hypothesis testing to localize root causes in multivariate time-series data. (8) CORAL [wang2023incremental]: A VAR-based online RCA method that decouples state-invariant and state-dependent information.
A-E Log Feature/Indicator Extraction
In this subsection, we provide the details of converting the raw data into the time series, though this is not within the scope of this work. Specifically, we first use the Drain parser [he2017drain] to transform the unstructured log event into structured log templates for each entity. Then, we partition the log data with fixed time windows, such as 5 minutes, and set time steps at 10 seconds. Within these intervals, we count the occurrence of each log template to derive the log frequency feature. The extraction of the log frequency feature is inspired by the insight that the recurrence of a log event template often correlates with its significance. For instance, when a microservice system experiences Distributed Denial of Service (DDoS) attacks, the system will produce an unusual volume of system logs, indicating abnormal activity. Thus, the log frequency provides the information to identify unusual patterns indicative of potential failure scenarios. In addition to log frequency, we also extract a second type of log feature known as the ‘golden signal.’ Notice that different from log frequency, golden signal heavily relies on domain knowledge and it only focuses on the abnormal system logs. More specifically, we are only interested in some keywords, including ‘error,’ ‘exception,’ ‘critical,’ ‘fatal’, and various others indicative of system anomalies. By identifying these keywords within log event templates, we can discern abnormal occurrences for system failure localization. Similar to the frequency-based feature, we compute the number of abnormal log events to derive the golden signal-based feature.
TABLE VII: Ablation Study for different neural network architecture
@@ A-E Log Feature/Indicator Extraction
TABLE VIII: Experimental results with more modality
Product Review Dataset PR@1 PR@5 PR@10 MRR MAP@3 MAP@5 MAP@10 Time (s)
OCEAN 1 1 1 1 1 1 1 20.1
OCEAN + trace 1 1 1 1 1 1 1 26.3
Train Ticket PR@1 PR@5 PR@10 MRR MAP@3 MAP@5 MAP@10 Time (s)
OCEAN 0.2 0.4 1 0.38 0.33 0.36 0.58 3.2
OCEAN + trace 0.2 0.6 1 0.39 0.33 0.44 0.62 3.8
A-F Experimental Results with More Modalities
OCEAN can naturally extend to include additional modalities, such as traces. These types of data can enhance the model’s ability to capture complex interactions and dependencies within the system. We conducted additional experiments by incorporating traces into the AIOps and Train Ticket datasets. The results in Table VIII demonstrated improved performance, as the inclusion of traces provided valuable context and enriched the causal structure learning. This additional information allows OCEAN to more accurately identify root causes and improve the precision of its analysis.
TABLE IX: Comparison with physical graph
Graphs SHD AUROC
Causal graph Learned From Metric Data 0.314 0.865
Causal graph Learned From Log Data 0.593 0.663
Fused Causal Graph 0.298 0.881
A-G Comparison with Physical Graph
Here, we evaluate the quality of the learned causal graph by comparing it with the physical dependency graph with two settings. In the first setting, we compared the causal graph learned by each modality (corresponding to the inter-modal graphs) and in the second setting, we compared the fused causal graph from two modality (corresponding to the intra-model graph). Following Dynotear [DBLP:conf/aistats/PamfilSDPGBA20], we use AUROC and SHD as two metrics to quantify the difference between learned causal graphs and the physical dependency graph. The experimental results are shown in Table IX.
Refer to caption Refer to caption Refer to caption
Figure 3: Parameter analysis on the Product Review dataset w.r.t MRR.
@@ A-G Comparison with Physical Graph
A-H Comparison with different architecture
We evaluate the efficiency of our long-term temporal causal relationship module by replacing the dilated convolutional neural network with LSTM or Transformer model on the Product Review and Train Ticket datasets in Table VII. The results showed a performance decline, demonstrating the effectiveness of our design.
A-I Parameter Analysis
@@
In this subsection, we present detailed parameter sensitivity analysis conducted on the Product Review, Online Boutique and Train Ticket datasets. we assess the impact of three parameters on the overall objective functions (Eq. (III-C)): $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$. The experimental results are displayed in Figures 3, 4 and 5, showing the Mean Reciprocal Rank (MRR) on the Product Review dataset. On these figures, the x-axis represents $\ln(\lambda_{i})$ for $i\in[1,2,3]$, and the y-axis shows the MRR score. Our analysis reveals that (1) We consistently observe that OCEAN favors a larger value of $\lambda_{1}$ on these two datasets as the temporal causal structure learning module is crucial in capturing both temporal and causal dependency. (2) Conversely, $\lambda_{2}$ and $\lambda_{3}$ exhibit optimal performance at relatively lower values (e.g., $\ln({\lambda_{2}})=-0.7$ and $\ln(\lambda_{3})=0$) on the Product Review dataset, with performance declining noticeably at higher levels. However, further reducing $\lambda_{2}$ and $\lambda_{3}$ also leads to diminished performance, which verifies the important role of sparse regularization and the acyclic constraint of the causal graph. Different from the parameter analysis on the Product Review dataset, we find that $\lambda_{2}$ and $\lambda_{3}$ are not very sensitive on the Online Boutique and Train Ticket datasets. We conjecture that this can be attributed to the small size of these two datasets and both sparse regularization and acyclic constraint contribute less to securing high performance on these two datasets than the Product Review dataset.
A-J Detailed Related Work
Root Cause Analysis. Current root cause analysis (RCA) methods can be categorized into two main branches: single-modal RCA methods and multi-modal RCA methods. Single-modal RCA methods primarily investigate causal relationships among system components using one type of data only [sporleder2019root, duan2020root, DBLP:conf/iwqos/MengZSZHZJWP20, soldani2022anomaly, aggarwal2020localization, li2021practical]. For instance, Liu et al. [DBLP:conf/icse/LiuH0LZGLOW21] generate a service call graph based on domain-specific software and rules, while Wang et al. [wang2023interdependent] construct causal networks from time series data. However, these methods often exhibit suboptimal performance due to their reliance on single-modal data. To enhance accuracy, recent research integrates multi-modal data for RCA [yu2023nezha, DBLP:conf/ispa/HouJWLH21, DBLP:conf/www/ZhengCHC24]. Nezha [yu2023nezha] and PDiagnose [DBLP:conf/ispa/HouJWLH21] extract and combine information from each modality individually, while MULAN [DBLP:conf/www/ZhengCHC24] and MM-DAG [lan2023mm] consider interactions among modalities, constructing comprehensive causal graphs. Despite notable progress, these approaches are implemented offline, necessitating extensive data collection and retraining for new faults. Wang et al. [wang2023incremental] enable online root cause identification by decoupling state-invariant and state-dependent information to learn a causal graph for root cause identification. However, their focus remains on single-modal data. Recently, large language model (LLM)-based approaches have emerged as a new research direction for learning causal relations in root cause identification, owing to the success of LLMs in tackling complex tasks [chen2024automatic, shan2024face, goel2024x, zhou2024causalkgpt, roy2024exploring, wang2023rcagent]. For example, Chen et al. [chen2024automatic] introduce RCACopilot, an on-call system powered by LLMs to automate RCA for cloud incidents. Similarly, Shan et al. [shan2024face] propose an approach that first identifies log messages indicating configuration-related errors, then localizes suspected root-cause configuration properties based on these log messages and LLM-generated configuration settings. While LLMs could effectively learn temporal dependencies within the metric data of individual system entities, they often struggle to capture interdependencies—such as causal relationships between different system entities—leading to higher computational costs. Unlike existing RCA methods, this paper addresses the online multi-modal RCA problem by uniquely modeling long-term temporal dependencies while simultaneously capturing the cross-modal correlation of multiple factors.